OpenStudy ($oumya):

what is the lowest value of2^sin^2x+2^cos^2x

OpenStudy (anonymous):

are you familiar with min/maxing functions using derivatives?

OpenStudy ($oumya):

no

OpenStudy (anonymous):

oh man...sorry then I can't help you.

ganeshie8 (ganeshie8):

Say, \(\large 2^{\sin^2x} + 2^{\cos^2x} = t\) is the min value \(\large 2^{\sin^2x} + 2^{1-\sin^2x} = t\) \(\large \left(2^{\sin^2x} \right)^2 + 2 = \left(2^{\sin^2x}\right) t\) plugin \(\large 2^{\sin^2x} = y\) \(\large y^2 -yt + 2 = 0\) Since \(y\) is always real (positive), D > 0 : \(\large t^2 - 4(1)(2) \ge 0 \) \(\large t \ge \sqrt{8} \)

OpenStudy (anonymous):

The function y = 2cos^2 x + 2 sin^2 x is always positive, Then, the lowest value of y is zero. We can prove by another way. y = 2 (sin^2 x + cos^2 x) = 2[(sin^2 x + cos^2 x)^2 - 2sin x*cos x) y is minimum when (2sin x*cos x = sin 2x) is max, meaning sin 2x = 1, then y = 0. It occurs when 2x = Pi/2 -> x = Pi/4.